40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 20 16:14:17 2017, gautam, Update, Equipment loan, Impedance test kit borrowed from Downs 
    Reply  Thu Sep 21 01:32:00 2017, gautam, Update, ALS, AUX X Innolight AM measurement running Innolight_AM.pdfInnolight_AM.tar.gzIMG_7599.JPG20170921_203741_TFAG4395A_21-09-2017_115547_FourSquare.pdf
       Reply  Thu Sep 21 01:55:16 2017, rana, Update, ALS, X End table of Shame 20170921_003143.jpg20170921_002430.jpg20170921_002243.jpg20170921_001906.jpg
          Reply  Thu Sep 21 15:23:04 2017, gautam, Omnistructure, ALS, Long cable from LSC->IOO 
          Reply  Fri Sep 29 11:16:52 2017, Steve, Update, ALS, Y End table corrected ETMYf1.jpg
          Reply  Fri Oct 6 17:08:09 2017, Steve, Update, ALS, X End table beam traps corrected c1.jpgc2.jpgc3.jpgc4.jpg
       Reply  Sun Sep 24 20:47:15 2017, rana, Update, Computer Scripts / Programs, RF TF Uncertainties TFAG4395A_21-09-2017_115547_FourSquare.pdfTFAG4395A_21-09-2017_115547.tgz
          Reply  Wed Sep 27 00:20:19 2017, gautam, Update, ALS, More AM sweeps TFAG4395A_26-09-2017_202344_FourSquare.pdflowTemp.tgzTFAG4395A_26-09-2017_231630_FourSquare.pdfhighTemp.tgz
             Reply  Wed Sep 27 23:44:45 2017, gautam, Update, ALS, Proposed PM measurement setup IMG_7609.JPG
Message ID: 13335     Entry time: Wed Sep 27 00:20:19 2017     In reply to: 13329     Reply to this: 13337
Author: gautam 
Type: Update 
Category: ALS 
Subject: More AM sweeps 

Attachment #1: Result of AM sweeps with EX laser crystal at nominal operating temperature ~ 31.75 C.

Attachment #2: Tarball of data for Attachment #1.

Attachment #3: Result of AM sweeps with EX laser crystal at higher operating temperature ~ 40.95 C.

Attachment #4: Tarball of data for Attachment #2.


Remarks:

  • Confirmed that PDA 55 is in the "0dB" setting - the actual dial is unmarked, and has 5 states. I guessed that the left-most one is 0dB, and checked that if I twiddled the dial by one state to the right, the DC level on the scope increased by 10dB as advertized. Didn't check all the states.
  • DC level is ~2.3V on a high-impedance scope. So it will be ~1.15V to a 50ohm load, which is what the DC block is. The inverse of this value is used to calibrate the vertical axis of the TF measurement to RIN/V.
  • Input R (split RF source signal) attenuation: 20dB. Input A (PDA55 output) attenuation: 0dB.
  • Main problem is still network hangups when trying to do many sweeps.
  • Seems to persist even when I connect the GPIB box to one of the network switches - so don't think we can blame the WiFi.
  • Need to explore possibility of speedup - takes >2hours to run ~50scans!

To-do:

  • Overlay median and uncertainty plots for the two temp. settings. There is a visible diference in both the locations and depths/heights of various notches/peaks in the AM profile.
  • Repeat test with a fast focusing lens to focus the beam more tightly on the PD active area to confirm that the measured AM is indeed due to the PZT drive and not from beam-jitter (presently, spot diameter is ~0.5x active area diameter, to eye).
  • Get the PM data.
  • Depending on what the PM data looks like, do a more fine-grained scan around some promising AM notches / PM peaks.
Attachment 1: TFAG4395A_26-09-2017_202344_FourSquare.pdf  2.523 MB  | Hide | Hide all
TFAG4395A_26-09-2017_202344_FourSquare.pdf
Attachment 2: lowTemp.tgz  406 kB  Uploaded Wed Sep 27 01:24:52 2017
Attachment 3: TFAG4395A_26-09-2017_231630_FourSquare.pdf  2.260 MB  Uploaded Wed Sep 27 01:25:08 2017  | Hide | Hide all
TFAG4395A_26-09-2017_231630_FourSquare.pdf
Attachment 4: highTemp.tgz  371 kB  Uploaded Wed Sep 27 01:25:25 2017
ELOG V3.1.3-